OpenStudy (anonymous):

The gravitational force exerted on a baseball is 2.28 N down. A pitcher throws the ball horizontally with velocity 14.0 m/s by uniformly accelerating it along a straight horizontal line for a time interval of 182 ms. The ball starts from rest.

OpenStudy (anonymous):

An electron of mass 9.11 10-31 kg has an initial speed of 2.20 105 m/s. It travels in a straight line, and its speed increases to 6.20 105 m/s in a distance of 4.20 cm. Assume its acceleration is constant. (a) Determine the magnitude of the force exerted on the electron.

OpenStudy (anonymous):

\[v^2=u^2+2as\]\[F=ma\]Therefore\[F=\frac{m(v^2-u^2)}{2s}\]Where v = final velocity u = initial velocity a = acceleration s = displacement F = force m = mass Now it's a simple matter of feeding numbers into a calculator!